OpenStudy (anonymous):

I need help matching Arithmetic and Geometric ratios? Graph is included! MEDAL

OpenStudy (anonymous):

 I am suppose to look at the term and choose the correct definition that goes with it. Cant figure it out. Please Help

OpenStudy (imstuck):

Did you get this or do you still need help?

OpenStudy (anonymous):

i still need help @IMStuck

OpenStudy (anonymous):

the first one is GP of common ratio 1/4

OpenStudy (anonymous):

like 128*1/4=32*1/4=8*1/4=2

OpenStudy (anonymous):

do you know the others

OpenStudy (anonymous):

second one is GP of ratio 3

OpenStudy (anonymous):

1*3=3*3=9*3=27

OpenStudy (anonymous):

3rd one is AP of common difference of 5

OpenStudy (anonymous):

4th one is again an AP of common difference -3

OpenStudy (anonymous):

any doubt? please ask